OpenStudy (anonymous):

how to parameterize x^3 + y^3 = 1?

OpenStudy (anonymous):

Can you solve for y in terms of x?

OpenStudy (anonymous):

z=0 by the way

OpenStudy (amistre64):

x = cos^3(t) y= sin^3(t) z = 0 right?

OpenStudy (anonymous):

the answer is given as the same as what you put, but the powers are (2/3), not 3 and i'm unsure how they got that

OpenStudy (amistre64):

I was close then :) I dont know how they got ^(2/3) either... id have to go back many years to remember ;)

OpenStudy (amistre64):

its a cubic equation that crosses the x and y axis 3 times.... i see that much

OpenStudy (anonymous):

i tried to use http://tutorial.math.lamar.edu/Classes/CalcIII/LineIntegralsPtI.aspx to help but :( cant seem to figure it out

OpenStudy (amistre64):

lets see if I can raeson this..... or muddy the waters :) x = rcos ; y = rsin ; 1 = sin^2 + cos^2 right?

OpenStudy (amistre64):

r^3 (cos^3 + sin^3) = sin^2 + cos^2

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

I see where the ^1/3 is coming from.... just how to get it :)

OpenStudy (anonymous):

it seems as though thats not too far away

OpenStudy (anonymous):

Do you see that if you plug in x = (cos theta)^{2/3) and y=(sin theta)^(2/3) you get the original equation? That's why the answer is correct, but doesn't give much insight into how you were supposed to guess it.

OpenStudy (amistre64):

yeah, the solution being correct is the easy part ;)

OpenStudy (amistre64):

1 is also the "radius" or vector measure from the origin.... we could try that route.

OpenStudy (amistre64):

are there any trig identitied we could utilize for this?

OpenStudy (anonymous):

ermm, not that i can think of

OpenStudy (amistre64):

r^3 (cos^3 + sin^3) = r^2 ; r^2 = 1 so r^3 = 1 ?

OpenStudy (anonymous):

ye, i think r^3 must have to be 1

OpenStudy (amistre64):

cos^3 + sin^3 = cos^2 + sin^2

OpenStudy (amistre64):

cos^3 - cos^2 = sin^3 - sin^3 ??

OpenStudy (amistre64):

typoed it...but you can see that lol

OpenStudy (anonymous):

ye, i see lol...what do you have to do to a square power to make it ^2/3? do you cube root it?

OpenStudy (amistre64):

yes; ^2/3 means cbrt(^2)

OpenStudy (anonymous):

it's because you want cosx^2+sinx^2=1. So if x=cost^(2/3) when you plug it in, the square will distrbute in to leave cos^2

OpenStudy (anonymous):

ok, i think i get it, thanks so much for the help!